OpenStudy (anonymous):

Suppose that the time it takes to do a job is inversely proportional to the number of workers. That is, the more workers on the job the less time required to complete the job. If it takes 3 workers 16 days to finish a job, how long will it take 8 workers? 6 days 10 days 2 days 14 days

OpenStudy (anonymous):

time = k (constant of proportionality) x no.of workers 16 days = k x 3 k= 16/3 time = k x 8 time = 16/3 x 8